OpenStudy (josh0404):

sketch the graph of piece-wise function given by y=x+1 if x0 and find the domain and range please help me guys?

jimthompson5910 (jim_thompson5910):

do you know how to graph y = x+1 without worrying about restrictions?

OpenStudy (josh0404):

what do you mean about worrying about restriction sir?

jimthompson5910 (jim_thompson5910):

do you know how to graph y = x+1 on a coordinate grid?

OpenStudy (josh0404):

yes sir

OpenStudy (josh0404):

using vertex

jimthompson5910 (jim_thompson5910):

this is a linear equation. There is no vertex

OpenStudy (josh0404):

using (h,k) formula

jimthompson5910 (jim_thompson5910):

focus on y = x+1 for now. This is linear. There is no vertex

OpenStudy (josh0404):

the graph pf x+1 is x=0 and y=1 I am right or wrong?

jimthompson5910 (jim_thompson5910):

(0,1) is one point on this line. Yes

jimthompson5910 (jim_thompson5910):

what's another point?

OpenStudy (josh0404):

x^2+2 where x=0 and y=2

jimthompson5910 (jim_thompson5910):

if you plugged x = 1 into y = x+1, what is the value of y?

OpenStudy (josh0404):

y=2

jimthompson5910 (jim_thompson5910):

when x = 1, y = 2 put those together and (1,2) is another point on the graph of y = x+1

jimthompson5910 (jim_thompson5910):

plot (0,1) and (1,2). Draw a straight line through them The line would normally go off forever in both directions; however, there's a restriction on x it says that y = x+1 if x <= 0 so y = x+1 is only graphed if x is 0 or smaller. Otherwise, you don't graph

OpenStudy (josh0404):

i want to see the final graph would you please draw it?

jimthompson5910 (jim_thompson5910):

Here's y = x+1 normally. See attached

jimthompson5910 (jim_thompson5910):

Here is y = x+1 when you apply the restriction see attached notice how there is a closed circle at the endpoint (to indicate we include this point)

OpenStudy (josh0404):

here what i did before but my teacher wrong me.

OpenStudy (josh0404):

how about the complete graph sir i want to see?

OpenStudy (josh0404):

help me please :( ? i want to learn

jimthompson5910 (jim_thompson5910):

this is what the full piecewise function would look like. See attached

jimthompson5910 (jim_thompson5910):

let me know if that helps or not

OpenStudy (josh0404):

now i know why im wrong because i shade the blue one instead of open circle :D

OpenStudy (josh0404):

how about the domain and range sir?

jimthompson5910 (jim_thompson5910):

hopefully you can see how the domain is the set of all real numbers? Notice how we can pick any x value and it leads to a point on the graph

OpenStudy (josh0404):

how about the range sir is a set of real number but not equal to 2?

jimthompson5910 (jim_thompson5910):

we can do almost every real number for the range, but there's a gap from y = 1 to y = 2

jimthompson5910 (jim_thompson5910):

we can say that the range is the set of y values such that \(\Large y \le 1\) or \(\Large y > 2\)

OpenStudy (josh0404):